OpenStudy (anonymous):

18y^5-8y^4/2y^2 simplify as much as possible

OpenStudy (anonymous):

i got 9y^3+4y^2

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we can simplify the second term, using the rules of division of powers with the same basis: \[\frac{{8{y^4}}}{{2{y^2}}} = \frac{8}{2} \cdot {y^{4 - 2}} = ...?\] please simplify

OpenStudy (michele_laino):

sorry is your expression like this: \[\frac{{18{y^5} - 8{y^4}}}{{2{y^2}}} = ...?\]

OpenStudy (michele_laino):

so the first term is correct! \(9y^3\)

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

the second term is: \[\frac{{ - 8{y^4}}}{{2{y^2}}} = \frac{{ - 8}}{2} \cdot {y^{4 - 2}} = ...?\]

OpenStudy (anonymous):

-4y^2

OpenStudy (michele_laino):

right so your second term is: \(-4y^2\), and the complete result is: \(9y^3-4y^2\)

OpenStudy (anonymous):

thank you. can you check one more

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

x^2+10x+26/x+6 i have to get the quotient and the remainder. I got x+4 & 2 for the remainder

OpenStudy (michele_laino):

your answer is correct, since we have: \[\left[ {\left( {x + 6} \right) \cdot \left( {x + 4} \right)} \right] + 2 = {x^2} + 10x + 26\]

OpenStudy (anonymous):

thank you. sorry but can you check another one?

OpenStudy (michele_laino):

more explanation: it is a proof, since we have: (quotient times divisor) plus remainder = dividend

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

3v^2(v+8)-7(v+8) rewrite by factoring out (v+8) I got (v+8)(3v^2-7)

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thanks